On January 2, 2018 Smith Electronics purchased a machine to help automate operations. The Company that manufactured the machine had a standard price for this item at $25,000, but Smith was able to negotiate a 20% discount from standard price. However, as a result of the lower price Smith paid $1,200 of freight costs, hired an installation team for $1,300 to install the new machine. In addition, Smith had to rent a crane for $800 as part of the installation. The city required Smith to pay a one-time inspection fee of $700. Last, Smith incurred $1,000 training employees to use the new equipment. The equipment has a five year useful life with no salvage value.

1. For 2018 compute depreciation expense for straight-line and double declining balance.

2. Assume the equipment performs poorly and you decide to sell it after 9 months. You are able to sell the equipment for $14,000 cash. You used double declining balance depreciation method.

* Record the sale
Answer #1
Cost of machine:
Machine standard price 25000
discount @20% 5000
Discounted price of Machine 20000
Freight costs 1200
Installation costs (1300+800) 2100
One time inspection fee 700
Cost of machine 24000
Requirement 1
Depreciation expense for 2018 using straight line method
= 24,000/5 = $4,800
Depreciation expense for 2018 using double declining balance method
= 2 X (24,000/5) = $9,600
Requirement 2
Depreciation expense for 9 months using double declining balance method
= 2 X (24,000/5) X 9/12 = $7,200
Machine value after 9 months
= $24000 -7,200 = $16,800
Loss on sale of machine = $16800 -$14,000 = $2,800
Entry for sale of Machine
Cash Account Debit                    $14,000
Loss on sale of Machine Debit   $2,800
Accumulated depreciation Debit $7,200
Machine credit                                             $24,000